OpenStudy (anonymous):

tan^2x-cot^2x=?? ill draw a pic

OpenStudy (anonymous):



OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{\sin^2x}{\cos^2x}-\frac{\cos^2x}{\sin^2x} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{\sin^2x\sin^2x}{\cos^2x\sin^2x}-\frac{\cos^2x\cos^2x}{\sin^2x\cos^2x} }\)

OpenStudy (anonymous):

okay so common denominators and i just cancel what needs to be canceled?

OpenStudy (solomonzelman):

you are right about the common denominator, but don't be too quick... tell me first, what do you get after adding the fractions.

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{\sin^2x\sin^2x}{\cos^2x\sin^2x}-\frac{\cos^2x\cos^2x}{\sin^2x\cos^2x} }\) \(\LARGE\color{black}{ \frac{\sin^2x\sin^2x-\cos^2x\cos^2x}{\cos^2x\sin^2x}}\)

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{\sin^2x(1-\cos^2x)-\cos^2x(1-\sin^2x)}{\cos^2x\sin^2x}}\) \(\LARGE\color{black}{ \frac{\sin^2x-\sin^2x\cos^2x~~-~~\cos^2x-\sin^2x\cos^2x}{\cos^2x\sin^2x}}\)

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{\sin^2x-2\sin^2x\cos^2x-\cos^2x}{\cos^2x\sin^2x}}\)

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{\sin^2x-2\sin^2x\cos^2x-(1-\sin^2x)}{\cos^2x\sin^2x}}\) \(\LARGE\color{black}{ \frac{\sin^2x-2\sin^2x\cos^2x-1+\sin^2x}{\cos^2x\sin^2x}}\)

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{2\sin^2x-2\sin^2x\cos^2x-1}{\cos^2x\sin^2x}}\)

OpenStudy (solomonzelman):

then just re-write each term.

OpenStudy (anonymous):

okay got it thanks :)

OpenStudy (solomonzelman):

just in case,